# SMoFi: Step-wise Momentum Fusion for Split Federated Learning on Heterogeneous Data

### v1

@@ 1 Introduction
The proliferation of mobile and sensing devices (Gubbi et al. 2013; Fortino and Trunfio 2014) has resulted in rich data at the edge, enabling a new range of Artificial Intelligence of Things applications (Wang et al. 2020b; Verbraeken et al. 2020). In this context, Federated Learning (FL) has been proposed as an important learning paradigm that exploits data generated at distributed edge devices while preserving data privacy (Konečnỳ et al. 2016; McMahan et al. 2017a; Bonawitz et al. 2019). Training large models on edge devices, however, remains a key challenge due to the limited computing resources at the edge. To expedite on-device model training, prior work has explored collaborative model training, resorting to richer computing resources  (Li et al. 2018; Eshratifar, Abrishami, and Pedram 2019; Wang et al. 2021). A promising approach is split learning (Gupta and Raskar 2018; Vepakomma et al. 2018b; Singh et al. 2019; Gao et al. 2020; Thapa et al. 2022), which divides a model into submodels trained separately on the edge and cloud, thereby offloading partial training overhead to the powerful server.
The effectiveness of split FL, in terms of both model performance and convergence, is largely undermined by data heterogeneity–i.e., non-IID (non-identically and/or independently distributed) data silos–often presenting in real-world scenarios. Training on such non-IID data introduces inconsistency in model updates across both client-side and server-side submodels; such a divergence accumulates as the training proceeds over iterations. Consequently, aggregation of diverse model updates results in a slower convergence and inferior performance of the global model compared with the IID setting. Existing methods to address such a challenge in FL either modify the loss function (Li et al. 2020; Li, He, and Song 2021; Gao et al. 2022) or impose robustness constraints on server aggregation (Hsu, Qi, and Brown 2019; Wang et al. 2020a; Reddi et al. 2020; Shi et al. 2025). While many of them are adaptable to split FL, they overlook a unique characteristic of split learning: the server directly controls the learning pace of multiple surrogate server-side models (each paired with a client-side submodel) that typically constitute the majority of the full model. This leads us to ask: Can we impose constraints on model training in split FL by leveraging its inherent client-server interaction–enforced more tightly and synchronously than in conventional FL–without introducing additional overheads or privacy risk?
@@
  • We propose a novel split FL framework that improves the consistency of server-side submodels during learning on non-IID data through step-wise momentum fusion.
@@ 2.2 SMoFiDesign
Momentum Alignment. In the Parallel updating framework, the server maintains $\lvert\mathcal{J}^{n}\lvert$ optimizers where each optimizer has a momentum buffer $m_{s,j}^{(n,\tau)}$ recording the accumulated gradients of the $j$-th server-side model. The objective inconsistency causes a divergence between momentum buffers $\{m_{s,j}^{(n,\tau)}\}_{j\in\mathcal{J}^{n}}$, which accumulates as training progresses over iterations. The momentum buffers then further negatively affect gradient descent, causing the models across clients to converge in increasingly divergent directions. To address this, SMoFi aligns the momentum buffers across the server-side optimizers, and Equation 2 is then reformulated as
@@
where $\bar{m}_{s}^{(n,\tau)}$ is the aligned momentum buffer, synchronized after all server-side solvers perform gradient descent at each local step. Specifically, each server-side model is updated by the gradient of a unified momentum term combined with the stochastic gradients over local mini-batch samples (row 7 in Algorithm 1).
@@
Staleness Factor. To maintain the effectiveness of constraint throughout all $max\{T_{j}\}_{j\in\mathcal{J}^{n}}$ local steps, SMoFi introduces a mechanism that utilizes $\mathcal{H}^{(n)}$, a record of the momentum buffers $m_{s,j}^{(n,\tau+1)}$ when the current step reaches the final step of the client, i.e., when $\tau=T_{j}$, $\forall j\in\mathcal{J}^{(n,\tau)}$. At each local step, SMoFi aligns the momentum buffer for the next step (row 12 in Algorithm 1) by averaging momentums from both the current and historical (recorded in $\mathcal{H}^{(n)}$) state of optimizers:
@@ 17 $\bar{m}_{s}^{(n,\tau+1)}\leftarrow$ Updating by Equation 5
The full workflow of split FL integrating SMoFi is detailed in Appendix C.1. The central server [^2] In SFLV1/SFLV2, a fed server is used to aggregate the client-side submodels for the client-side global model. For simplicity, we assume that the central server performs the operations of the fed server. first aggregates both client-side and server-side submodels at the end of the communication round: $\bar{\mathcal{W}}^{n}=\sum_{j\in\mathcal{J}^{n}}p_{j}\mathcal{W}_{j}^{n}$ where $\mathcal{W}_{j}^{n}=[\mathcal{W}_{c,j}^{(n,T_{j})},\mathcal{W}_{s,j}^{(n,T_{j})}]$. Similar to work (Hsu, Qi, and Brown 2019), we update the global model with momentum: $\mathcal{W}^{n}=\mathcal{W}^{n-1}-m_{g}^{n}$. The global momentum buffer is updated following $m_{g}^{n}=\beta_{g}m_{g}^{n-1}+\mathcal{W}^{n-1}-\bar{\mathcal{W}}^{n}$ and $\beta_{g}$ is the global momentum coefficient.
@@ 4.1 Experimental Setups
Baselines. The baselines chosen for comparison are categorized into momentum-based methods and split FL methods. We compare SMoFi against momentum-based counterparts on three baselines: 1) the vanilla FL framework FedAvg (McMahan et al. 2017b); 2) FedProx (Li et al. 2020), adding a proximal term into the local objective function; and 3) FedNAR (Li et al. 2023) with self-adjusted weight decay. Building upon each baseline, we integrate three momentum-based methods including: 1) FedAvgM (Hsu, Qi, and Brown 2019), applying server momentum during global updates; 2) SlowMo (Wang et al. 2019b), periodically synchronizing and updating the local momentum across clients; and 3) FedNAG (Yang et al. 2022), implementing Nesterov Accelerated Gradient (NAG) (Sutskever et al. 2013; Bengio, Boulanger-Lewandowski, and Pascanu 2013) for local training with periodic momentum aggregation. Additionally, we include three split FL frameworks for evaluation: SFLV1, SFLV2 (Thapa et al. 2022), and MergeSFL (Liao et al. 2024). For SFLV1, we investigate two server-side aggregation frequencies: after each step ($\bar{\tau}=1$) and after each local epoch ($\bar{\tau}=E$), and SFLV1 is equivalent to FedAvg when aggregating surrogate server-side models at each communication round.
@@ D.1 Preliminaries
$\mathbb{E}[\|\hat{\mathcal{W}}^{n}-\hat{\mathcal{W}}^{*}\|^{2}_{2}]$ (21)
@@
$\hskip 20.00003pt+\mathbb{E}\big[\|\eta^{n}\!\sum_{\tau\in[T]}\nabla\mathcal{L}(\hat{\mathcal{W}}^{(n-1,\tau)})-\eta^{n}\!\sum_{j\in\mathcal{J}}p_{j}\!\sum_{\tau\in[T]}g_{j}^{(n,\tau)}\|^{2}_{2}\big],$
@@
$-2\eta^{n}\mathbb{E}[\langle\hat{\mathcal{W}}^{n-1}-\hat{\mathcal{W}}^{*},\sum_{j\in{\mathcal{J}}}\sum_{\tau\in[T]}p_{j}\nabla\mathcal{L}_{j}(\hat{\mathcal{W}}^{(n-1,\tau)})\rangle]$ (23)
@@
[…]
$\hskip 20.00003pt+(\sigma^{2}+G^{2})(\eta^{n})^{2}T^{2}\lvert\mathcal{J}\lvert\sum_{j\in{\mathcal{J}}}p_{j}^{2}.$
@@
$\mathbb{E}[\|\eta^{n}\sum_{\tau\in[T]}\nabla\mathcal{L}(\hat{\mathcal{W}}^{(n-1,\tau)})-\eta^{n}\sum_{j\in\mathcal{J}}p_{j}\sum_{\tau\in[T]}g_{j}^{(n,\tau)}\|^{2}_{2}]$ (26)
@@
$=(\eta^{n})^{2}T^{2}\sigma^{2}\lvert\mathcal{J}\lvert\sum_{j\in\mathcal{J}}p_{j}^{2},$
@@
$\hskip 25.00003pt\leq\frac{v}{\gamma+n+1},$
@@
$\mathbb{E}[\|\hat{\mathcal{W}}^{n}-\Omega^{n}\|^{2}_{2}]$ $=\mathbb{E}[\|\hat{\mathcal{W}}^{n}-\hat{\mathcal{W}}^{n-1}+\hat{\mathcal{W}}^{n-1}-\Omega^{n}\|^{2}_{2}]$ (33)
@@
$=\frac{1}{\theta}(\eta^{n})^{2}G^{2}\lvert\mathcal{J}\lvert T^{2}\sum_{j\in\mathcal{J}}p_{j}^{2},$
@@
$\mathbb{E}[\mathcal{L}(\mathcal{W}^{N})]-\mathcal{L}(\mathcal{W}^{*})$ (36)
@@
$\leq\mathcal{O}(\frac{A}{(\gamma+N)})+\mathcal{O}(\frac{B}{(\gamma+N)})+\mathcal{O}(\frac{C}{{(\gamma+N)}}).$

### v2

@@ 1 Introduction
The proliferation of mobile and sensing devices (Gubbi et al. 2013; Fortino and Trunfio 2014) has resulted in rich data at the edge, enabling a new range of Artificial Intelligence of Things applications (Wang et al. 2020b; Verbraeken et al. 2020). In this context, Federated Learning (FL) has been proposed as an important learning paradigm that exploits data generated at distributed edge devices while preserving data privacy (Konečnỳ et al. 2016; McMahan et al. 2017a; Bonawitz et al. 2019). On-device training of large models, however, remains a key challenge due to limited computing resources at the edge. To expedite on-device model training, prior work has explored collaborative model training, resorting to richer computing resources  (Li et al. 2018; Eshratifar, Abrishami, and Pedram 2019; Wang et al. 2021). A promising approach is split learning (Gupta and Raskar 2018; Vepakomma et al. 2018b; Singh et al. 2019; Gao et al. 2020; Thapa et al. 2022), which divides a model into submodels trained separately on the edge and cloud, thereby offloading partial training overhead to the powerful server.
The effectiveness of split FL, in terms of both model accuracy and convergence, is largely undermined by data heterogeneity–i.e., non-IID (non-identically and/or independently distributed) data silos–often presenting in real-world scenarios. Training on such non-IID data introduces inconsistency in model updates across both client-side and server-side submodels; such a divergence accumulates as the training proceeds over iterations. Consequently, aggregation of diverse model updates results in inferior accuracy and slower convergence of the global model compared with the IID setting. Existing methods to address such a challenge in FL either modify the loss function (Li et al. 2020; Li, He, and Song 2021; Gao et al. 2022) or impose robustness constraints on server aggregation (Hsu, Qi, and Brown 2019; Wang et al. 2020a; Reddi et al. 2020; Shi et al. 2025). While many of them are adaptable to split context, they overlook a unique attribute of split learning: the server directly controls the learning pace of multiple surrogate server-side models (each paired with a client-side submodel) that typically constitute the majority of the full model. This leads us to ask: Can we impose constraints on model training in split FL by leveraging its inherent client-server interaction–enforced more tightly and synchronously than in conventional FL–without introducing additional overheads or privacy risk?
@@
  • We propose a novel split FL framework that improves the consistency of server-side submodel updates on non-IID data via step-wise momentum fusion.
@@ 2.2 SMoFiDesign
Momentum Alignment. In the Parallel updating framework, the server maintains $\lvert\mathcal{J}^{n}\lvert$ optimizers where each optimizer has a momentum buffer $m_{s,j}^{(n,\tau)}$ tracking the accumulated gradients of the $j$-th server-side model. Under objective inconsistency, these momentum buffers $\{m_{s,j}^{(n,\tau)}\}_{j\in\mathcal{J}^{n}}$ gradually diverge as training proceeds. The momentum buffers then further negatively affect gradient descent, causing the models across clients to converge in increasingly divergent directions. To address this, SMoFi aligns the momentum buffers across the server-side optimizers, and Equation 2 is then reformulated as
@@
where $\bar{m}_{s}^{(n,\tau)}$ is the aligned momentum, synchronized after all server-side solvers perform gradient descent at each local step. In this way, each server-side model is updated based on stochastic gradients over its local mini-batch samples, combined with a unified momentum term (line 7 in Algorithm 1).
@@
Staleness Factor. To maintain the effectiveness of constraint throughout all $max\{T_{j}\}_{j\in\mathcal{J}^{n}}$ local steps, SMoFi introduces a mechanism that leverages $\mathcal{H}^{(n)}$, a record of the momentum buffers $m_{s,j}^{(n,\tau+1)}$ at the final step of the client, i.e., when $\tau=T_{j}$, $\forall j\in\mathcal{J}^{(n,\tau)}$. At each local step, SMoFi aligns the momentum buffer for the next step (line 12 in Algorithm 1) by averaging both the current momentum of optimizers and the historical state stored in $\mathcal{H}^{(n)}$:
@@ 17 $\bar{m}_{s}^{(n,\tau+1)}\leftarrow$ Updating by Equation 5
The full workflow of split FL integrating SMoFi is detailed in Appendix C.1. The central server [^2] In SFLV1/SFLV2, a fed server is used to aggregate the client-side submodels for the client-side global model. For simplicity, we let the central server performs the operations of the fed server. first aggregates both client-side and server-side submodels at the end of the communication round: $\bar{\mathcal{W}}^{n}=\sum_{j\in\mathcal{J}^{n}}p_{j}\mathcal{W}_{j}^{n}$ where $\mathcal{W}_{j}^{n}=[\mathcal{W}_{c,j}^{(n,T_{j})},\mathcal{W}_{s,j}^{(n,T_{j})}]$. Similar to work (Hsu, Qi, and Brown 2019), we update the global model with momentum: $\mathcal{W}^{n}=\mathcal{W}^{n-1}-m_{g}^{n}$. The global momentum buffer is updated following $m_{g}^{n}=\beta_{g}m_{g}^{n-1}+\mathcal{W}^{n-1}-\bar{\mathcal{W}}^{n}$ and $\beta_{g}$ is the global momentum coefficient.
@@ 4.1 Experimental Setups
Baselines. The baselines chosen for comparison are categorized into momentum-based methods and split FL methods. We compare SMoFi against momentum-based counterparts on three baselines: 1) the vanilla FL framework FedAvg (McMahan et al. 2017b); 2) FedProx (Li et al. 2020), adding a proximal term into the local objective function; and 3) FedNAR (Li et al. 2023) with self-adjusted weight decay. Building upon each baseline, we integrate three momentum-based methods including: 1) FedAvgM (Hsu, Qi, and Brown 2019), applying server momentum during global updates; 2) SlowMo (Wang et al. 2019b), periodically synchronizing and updating the local momentum across clients; and 3) FedNAG (Yang et al. 2022), implementing Nesterov Accelerated Gradient (NAG) (Sutskever et al. 2013; Bengio, Boulanger-Lewandowski, and Pascanu 2013) for local training with periodic momentum aggregation. Additionally, we include three split FL frameworks for evaluation: SFLV1, SFLV2 (Thapa et al. 2022), and MergeSFL (Liao et al. 2024). For SFLV1, we investigate two server-side aggregation frequencies: after every training step ($\bar{\tau}=1$) and after each local epoch ($\bar{\tau}=E$). Note that the performance of SFLV1 is equivalent to FedAvg when aggregating surrogate server-side models at each communication round.
@@ D.1 Preliminaries
$\mathbb{E}[\|\hat{\mathcal{W}}^{n}-\hat{\mathcal{W}}^{*}\|^{2}_{2}]$
@@
$\hskip 20.00003pt+\mathbb{E}\big[\|\eta^{n}\!\sum_{\tau\in[T]}\nabla\mathcal{L}(\hat{\mathcal{W}}^{(n-1,\tau)})-\eta^{n}\!\sum_{j\in\mathcal{J}}p_{j}\!\sum_{\tau\in[T]}g_{j}^{(n,\tau)}\|^{2}_{2}\big],$ (21)
@@
$-2\eta^{n}\mathbb{E}[\langle\hat{\mathcal{W}}^{n-1}-\hat{\mathcal{W}}^{*},\sum_{j\in{\mathcal{J}}}\sum_{\tau\in[T]}p_{j}\nabla\mathcal{L}_{j}(\hat{\mathcal{W}}^{(n-1,\tau)})\rangle]$
@@
$\mathbb{E}[\|\hat{\mathcal{W}}^{n-1}-\hat{\mathcal{W}}^{*}-\eta^{n}\sum_{\tau\in[T]}\nabla\mathcal{L}(\hat{\mathcal{W}}^{(n-1,\tau)})\|^{2}_{2}]$
@@
$\hskip 20.00003pt+24(2\sigma^{2}+G^{2})(\eta^{n})^{3}T^{3}L$
$\hskip 20.00003pt+(\sigma^{2}+G^{2})(\eta^{n})^{2}T^{2}\lvert\mathcal{J}\lvert\sum_{j\in{\mathcal{J}}}p_{j}^{2}.$ (25)
@@
$\mathbb{E}[\|\eta^{n}\sum_{\tau\in[T]}\nabla\mathcal{L}(\hat{\mathcal{W}}^{(n-1,\tau)})-\eta^{n}\sum_{j\in\mathcal{J}}p_{j}\sum_{\tau\in[T]}g_{j}^{(n,\tau)}\|^{2}_{2}]$
@@
$=(\eta^{n})^{2}T^{2}\sigma^{2}\lvert\mathcal{J}\lvert\sum_{j\in\mathcal{J}}p_{j}^{2},$ (26)
@@
$\hskip 25.00003pt\leq\frac{v}{\gamma+n+1},$ (28)
@@
$\mathbb{E}[\|\hat{\mathcal{W}}^{n}-\Omega^{n}\|^{2}_{2}]$ $=\mathbb{E}[\|\hat{\mathcal{W}}^{n}-\hat{\mathcal{W}}^{n-1}+\hat{\mathcal{W}}^{n-1}-\Omega^{n}\|^{2}_{2}]$
@@
$=\frac{1}{\theta}(\eta^{n})^{2}G^{2}\lvert\mathcal{J}\lvert T^{2}\sum_{j\in\mathcal{J}}p_{j}^{2},$ (33)
@@
$\mathbb{E}[\mathcal{L}(\mathcal{W}^{N})]-\mathcal{L}(\mathcal{W}^{*})$
@@
$\leq\mathcal{O}(\frac{A}{(\gamma+N)})+\mathcal{O}(\frac{B}{(\gamma+N)})+\mathcal{O}(\frac{C}{{(\gamma+N)}}).$ (36)